# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients

### v1

@@ 4 Numerical Experiments
$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),$
$\phi^{1}(x)$ $=x^{2}(1-x)^{2},$
$\phi^{2}(y)$ $=y^{2}(1-y)^{2}.$
The variable advection coefficients are then given by:
$\sigma^{x}(x,y)=\sigma^{1}(x)\sigma^{2}(y),\qquad D_{y}(x,y)=\chi^{1}(x)\chi^{2}(y),$
where:
$\sigma^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma^{2}(y)=y^{2}(1-y)^{2},$
$\chi^{1}(x)$ $=x^{2}(1-x)^{2},\qquad\chi^{2}(y)=2y(1-3y+2y^{2}).$
We select the following rank-1 initial condition:
$f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|,\quad 0\leq x,y\leq 1.$
Since the solution will converge to a steady state up to a constant, we correct the mass of the initial condition to match that of the steady-state solution. To correct the mass of the solution, we adopt the LoMaC strategy outlined in [14, 13, 12], as implemented in [10]. Figure 8 shows the error of the numerical steady-state solution relative to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $19,050$ to $577,500$ as the mesh is refined, indicating the exceptionally large time steps allowed by the algorithm. The results confirm that second-order spatial convergence is achieved and demonstrate that, under these conditions, large time steps can be taken while still accurately reaching the correct steady state with asymptotic convergence.
Figure 8: Example 4.4. Log-log plot of error in the numerical steady-state solution compared to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, ranges from $19,050$ to $577,500$. The plot demonstrates the retrieval of second-order spatial convergence in the steady state.
@@ 5 Conclusion
We have proposed a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients. We discretize the PDE using second-order centered finite differences in space and implicitly in time with up to third-order accuracy using DIRK, resulting in a generalized Sylvester matrix equation. Our algorithm is based on several key strategies: constructing dimension-wise basis from extended Krylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation. This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{3})$, with $r$ being the rank of the solution during Krylov iteration process, effectively mitigating the curse of dimensionality and making it comparable in efficiency to constant-coefficient cases [10]. We have verified our algorithm through a series of numerical experiments. Overall, our adaptive-rank algorithm offers a scalable and efficient solution for time-dependent advection-diffusion type PDEs with variable coefficients. It effectively balances computational cost and accuracy by efficiently constructing and truncating the Krylov basis and dynamically adjusting the solution rank. The ability to handle large meshes and time steps without sacrificing accuracy makes it a promising tool for a wide range of applications. Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
@@ Acknowledgements
This work was partially supported by the Multifaceted Mathematics Integrated Capability Centers (MMICCs) program of DOE Office of Applied Scientific Computing Research (ASCR). Los Alamos National Laboratory (LANL) is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). J.Q. and H.E. were supported by NSF grant NSF-DMS-2111253. J.Q. was supported by Air Force Office of Scientific Research FA9550-22-1-0390, Department of Energy DE-SC0023164 and Air Force Office of Scientific Research (AFOSR) FA9550-24-1-0254 via the Multidisciplinary University Research Initiatives (MURI) Program. H.E. was also supported by MMICCs at LANL during the summer of 2024.
@@ Appendix A Discretization of the Diffusion Operator
We provide a detailed derivation of the matrix formulation of the discretization of the diffusion operator for the two-dimensional advection-diffusion equation with separable diffusion coefficients. We consider the discretization of the diffusion term in the $x$-direction, with other directions treated similarly. Consider the diffusion term in the $x$-direction:
@@
Our goal is to discretize this term using a 3-point stencil on a uniform grid with spacing $\Delta x$ in the $x$-direction. Let $u_{i,j}\approx u(x_{i},y_{j})$, with $i=0,\cdots,N_{x}-1$ and $j=0,\cdots,N_{y}-1$.
We approximate the spatial derivative using central differences. The discrete approximation at grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x}\left[\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i+\frac{1}{2},j}-\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i-\frac{1}{2},j}\right],$ (39)
where the fluxes at the cell faces are approximated by:
$\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i+\frac{1}{2},j}$ $\approx\phi^{x}_{i+\frac{1}{2},j}\cdot\frac{u_{i+1,j}-u_{i,j}}{\Delta x},$ (40)
$\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i-\frac{1}{2},j}$ $\approx\phi^{x}_{i-\frac{1}{2},j}\cdot\frac{u_{i,j}-u_{i-1,j}}{\Delta x}.$ (41)
Here $\phi^{x}_{i+\frac{1}{2},j}=\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ Substituting back into Eq. (39), we obtain:
$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x^{2}}\left[\phi^{1}_{i+\frac{1}{2}}(u_{i+1,j}-u_{i,j})-\phi^{1}_{i-\frac{1}{2}}(u_{i,j}-u_{i-1,j})\right]\phi^{2}_{j}.$ (42)
To express the discretization in matrix form, we define:
@@
We also define the forward and backward difference matrices in the $x$-direction for the interior nodes $i=1,\cdots,N_{x}-2$, where $i=0$ and $i=N_{x}-1$ correspond to zero Dirichlet boundary conditions and rest at the boundary:
$D_{+}$ $=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},$ (43)
$D_{-}$ $=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}.$ (44)
Using these definitions, the discrete diffusion term in Eq. (42) can be represented in the matrix form as:
@@
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Phi}_{xx}$ in Equation (6). The discretization in the y-direction follows the same strategy.
@@ Appendix B Discretization of the Advection Operator
We provide a detailed derivation of the matrix formulation of the discretization of the advection operator for the two-dimensional advection-diffusion equation with variable, separable advection coefficients. As before, we focus on the discretization of the advection term in the $x$-direction:
@@
Our goal is to discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction. Let $u_{i,j}\approx u(x_{i},y_{j})$. To preserve linearity of our discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative. This discretization is not monotonicity-preserving unless the so-called cell Reynolds number condition is satisfied:
$\text{Re}_{c}=\frac{\sigma_{\text{max}}\Delta x}{\phi_{\text{max}}}<2,$ (48)
where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity, and $\phi_{\text{max}}$ is the maximum diffusion coefficient. This condition is practical in our formulation due to the very favorable scaling of our low-rank representation with mesh size and rank.
The discrete approximation at grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}\approx\frac{1}{\Delta x}\left[\left(\sigma^{x}u\right)_{i+\frac{1}{2},j}-\left(\sigma^{x}u\right)_{i-\frac{1}{2},j}\right],$ (49)
where the fluxes at the cell faces are approximated by:
$\left(\sigma^{x}u\right)_{i+\frac{1}{2},j}$ $\approx\sigma^{1}_{i+\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i+1,j}+u_{i,j}}{2},$ (50)
$\left(\sigma^{x}u\right)_{i-\frac{1}{2},j}$ $\approx\sigma^{1}_{i-\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i,j}+u_{i-1,j}}{2}.$ (51)
Here, $\sigma^{1}_{i+\frac{1}{2}}=\sigma^{1}(x_{i+\frac{1}{2}})$ is the advection coefficient evaluated at the midpoint between $x_{i}$ and $x_{i+1}$, and similarly for $\sigma^{1}_{i-\frac{1}{2}}$. Substituting back into the discrete approximation, we have:
$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}$ $\approx\frac{1}{\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i+1,j}+u_{i,j}}{2}-\sigma^{1}_{i-\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i,j}+u_{i-1,j}}{2}\right]$ (52)
$=\frac{1}{2\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}(u_{i+1,j}+u_{i,j})-\sigma^{1}_{i-\frac{1}{2}}(u_{i,j}+u_{i-1,j})\right]\sigma^{2}_{j}.$ (53)
To express the discretization in matrix form, we define:
@@
$S_{+}$ $=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},$ (54)
$S_{-}$ $=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}.$ (55)
@@ C.1 DIRK Scheme
DIRK methods are characterized by the Butcher tableau, as shown in Table 1.
@@
We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage. For instance the Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
@@
Table 2: DIRK2 Butcher table with $\gamma=1-\frac{\sqrt{2}}{2}$.
@@
Table 3: DIRK3 Butcher table with $x=0.4358665215$.
Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (8) results in a generalized Sylvester equation that must be solved at each DIRK stage. The corresponding DIRK scheme for the matrix differential equation (8) from $t^{(0)}$ to $t^{(1)}$ is written as:
@@
More compactly, within each stage $k$ of the DIRK scheme, one must minimize the residual equation (32) given by:
${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$
As shown in Equation (33), the residual equations lead to a generalized Sylvester equation to be solved at each stage.
@@ C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$
Expanding this residual using $\tilde{\mathscr{L}}$, defined in (14), leads to the $k$th-stage Sylvester equation:
@@
We follow the strategy discussed in [10] to perform high-order integration. The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${S}_{1}$ through the stages. The proposed procedure is as follows:
@@
      Solve reduced Sylvester equation for $S^{(k)}$ using preconditioned GMRES,
@@
      where $\tilde{P}_{1}=U_{1}^{T}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{T}P_{1}V_{1}$, and is valid for all stages since the basis constructed is fixed throught the DIRK stages by assumption. Solve for $S^{(k)}$ and obtain the intermediate RK solutions as $U_{1}S^{(k)}V_{1}^{\top}$. To further improve computational efficiency, from (59), we have,
      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({S^{(\ell)}}-{\tilde{B}^{(\ell)}}\right),$
      leading to an efficient computation of $\tilde{B}^{(k)}$:
      $\tilde{B}^{(k)}=\tilde{B}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({S^{(\ell)}}-{\tilde{B}^{(\ell)}}\right),$ (61)
      with $\tilde{B}_{1}$ defined in (19). This avoids the need of evaluating the full size $B^{(k)}$ in (32).
@@
    Perform truncation of evolved solution, i.e. $\{U_{1},S_{1},V_{1}^{\top}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}^{\top}\})$, with $S_{1}:=S_{1}^{(s)}$.
@@
    $\left\|{{\bf R}^{(s)}_{\mathscr{L}}}\right\|=\left\|{R}_{U}{\text{diag}\left(S_{1}-\tilde{B}^{(s)}_{1},-a_{ss}\Delta t\left(I_{{\mathcal{R}}_{\text{ranks}}}\otimes S_{1}\right)\right)}{R}_{V}^{\top}\right\|,$
    where $R_{U}$ and $R_{V}$ are upper triangular matrices from a reduced QR decomposition as described in (23), which can be done once for all DIRK stages.
  5. 1.
    Compare the computed residual norm $\|{\bf R}^{(s)}_{\mathscr{L}}\|$ with the given error tolerance $\epsilon_{tol}$. If the tolerance is met, then admit the solution with ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$; otherwise, we go back to Step 1 to further augment Krylov subspaces.

### v2

@@ 4 Numerical Experiments
$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),\qquad\phi^{1}(x)=x^{2}(1-x)^{2},\qquad\phi^{2}(y)=y^{2}(1-y)^{2}.$
The variable advection coefficients are $\sigma^{x}(x,y)=\sigma_{1}^{1}(x)\sigma_{1}^{2}(y)$, $\sigma^{y}(x,y)=\sigma_{2}^{1}(x)\sigma_{2}^{2}(y),$ where:
$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\qquad\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\qquad\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
We select the following rank-1 initial condition: $f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|$. The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-3}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-8}$, the SVD truncation threshold is $\epsilon=10^{-8}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$.
Figure 8 shows the error of the numerical steady-state solution, reached using ten time steps, relative to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $1,190$ to $36,076$ as the mesh is refined. The results are obtained by rescaling the numerical solution so that its normalization constant is the same as the analytical one. They confirm second-order spatial convergence and demonstrate the ability of our algorithm to take large timesteps to reach the correct steady state with expected asymptotic accuracy.
Figure 8: Example 4.3. Log-log plot of error in the numerical steady-state solution compared to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, ranges from $1,190$ to $36,076$. The plot shows second-order spatial convergence in the steady state.
@@ 5 Conclusion
We have proposed a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients. Our algorithm is based on several key strategies: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation. This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality, effectively mitigating the curse of dimensionality and making it comparable in efficiency to the constant-coefficient case [14]. Such complexity scaling is numerically verified with our extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances. Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
@@ Acknowledgements
This work was partially supported by the Multifaceted Mathematics Integrated Capability Centers (MMICCs) program of DOE Office of Applied Scientific Computing Research (ASCR). Los Alamos National Laboratory (LANL) is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). J.Q. and H.E. were supported by NSF grant NSF-DMS-2111253. J.Q. was also supported by Air Force Office of Scientific Research FA9550-22-1-0390, Department of Energy DE-SC0023164 and Air Force Office of Scientific Research (AFOSR) FA9550-24-1-0254 via the Multidisciplinary University Research Initiatives (MURI) Program. H.E. was also supported by MMICCs at LANL during the summer of 2024.
@@ Appendix A Discretization of the Diffusion Operator
We derive the matrix formulation of the discretization of the diffusion operator for the two-dimensional advection-diffusion equation with separable diffusion coefficients. We consider the discretization of the diffusion term in the $x$-direction, with other directions treated similarly:
@@
We discretize this term using a 3-point stencil on a uniform grid with spacing $\Delta x$ in the $x$-direction. Let $u_{i,j}\approx u(x_{i},y_{j})$, with $i=0,\cdots,N_{x}-1$ and $j=0,\cdots,N_{y}-1$. We approximate the spatial derivative using central differences. The discrete approximation at grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x^{2}}\left[\phi^{1}_{i+\frac{1}{2}}(u_{i+1,j}-u_{i,j})-\phi^{1}_{i-\frac{1}{2}}(u_{i,j}-u_{i-1,j})\right]\phi^{2}_{j},$ (39)
with $\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ To express the discretization in matrix form, we define:
@@
We also define the forward and backward difference matrices in the $x$-direction for the interior nodes $i=1,\cdots,N_{x}-2$. The nodes $i=0$, $i=N_{x}-1$ rest at the boundary and enforce homogeneous Dirichlet boundary conditions (other boundary conditions can be readily considered):
$D_{+}=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},D_{-}=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}.$ (40)
Using these definitions, the discrete diffusion term in Eq. (39) can be represented in matrix form as:
@@
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Phi}_{xx}$ in (9). The discretization in the $y$-direction follows the same strategy.
@@ Appendix B Discretization of the Advection Operator
We derive the matrix formulation of the discretization of the advection operator for the two-dimensional advection-diffusion equation with variable, separable advection coefficients. As before, we focus on the discretization of the advection term in the $x$-direction:
@@
We discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction. To preserve linearity of our discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative. This discretization is monotonicity-preserving when the following condition is satisfied:
$\Delta x<2\frac{\phi_{\text{max}}}{\sigma_{\text{max}}},$ (44)
where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity and $\phi_{\text{max}}$ is the maximum diffusion coefficient. This condition is practical in our formulation due to the very favorable scaling of our low-rank representation with mesh size and rank.
The discrete approximation at the grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}$ $\approx\frac{1}{2\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}(u_{i+1,j}+u_{i,j})-\sigma^{1}_{i-\frac{1}{2}}(u_{i,j}+u_{i-1,j})\right]\sigma^{2}_{j},$ (45)
where $\sigma^{1}_{i+\frac{1}{2}}=\sigma^{1}(x_{i+\frac{1}{2}})$ is the advection coefficient evaluated at the midpoint between $x_{i}$ and $x_{i+1}$, and similarly for $\sigma^{1}_{i-\frac{1}{2}}$, and $\sigma^{2}_{j}$ is evaluated at $y_{j}$. To express the discretization in matrix form, we define:
@@
$S_{+}=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},S_{-}=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}.$ (46)
@@ C.1 DIRK Scheme
DIRK methods are characterized by the Butcher tableau, Table 1.
@@
We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage. The Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
Table 2: DIRK2 Butcher table, $\gamma=1-\frac{\sqrt{2}}{2}$.
@@
Table 3: DIRK3 Butcher table, $x=0.4358665215$.
@@
Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (11) results in a GSE that must be solved at each DIRK stage. The corresponding DIRK scheme for the matrix differential equation (11) from $t^{(0)}$ to $t^{(1)}$ is written as:
@@
Within each stage $k$ of the DIRK scheme, one must solve the residual equation given by:
Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage GSE:
For the DIRK update, we adopt the strategy outlined in [14], with minor modifications to improve efficiency. In particular, we: (i) construct the basis using the time-step of the first stage and retain it for subsequent stages, (ii) evolve the projected system for the $S_{1}$ matrix across the DIRK stages, and (iii) truncate and evaluate the residual to determine whether to admit the solution or return to step (i) to further augment the basis. Key modifications vs. [14] include truncating the solution before the residual computation, and only checking the residual at the end of the DIRK step. The reduced matrix is solved using preconditioned GMRES, as before.
@@ C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
Expanding this residual using $\tilde{\mathscr{L}}$, defined in (16), leads to the $k$th-stage SE:
@@
We follow [14] to perform high-order integration. The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${{\bf S}}_{1}$ through the stages. The proposed procedure is as follows:
@@
      Solve reduced SE for ${\bf S}^{(k)}$ using preconditioned GMRES,
@@
      where $\tilde{P}_{1}=U_{1}^{T}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{T}P_{1}V_{1}$, and is valid for all stages since the basis constructed is fixed throught the DIRK stages by assumption. Solve for ${\bf S}^{(k)}$ and obtain the intermediate RK solutions as $U_{1}S^{(k)}V_{1}^{\top}$. To further improve computational efficiency, from (53), we have,
      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$
      leading to an efficient computation of $\tilde{{\bf B}}^{(k)}$:
      $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$ (55)
      with $\tilde{{\bf B}}_{1}$ defined in (21). This avoids the need to evaluate the full size ${\bf B}^{(k)}$ in (51).
@@
    Perform truncation of evolved solution, i.e. $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}^{\top}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}^{\top}\})$, with $S_{1}:=S_{1}^{(s)}$.
@@
    $\left\|{{\bf R}^{(s)}_{\mathscr{L}}}\right\|=\left\|{R}_{U}{\text{diag}\left(-\tilde{B}^{(s)}_{1},S_{1,\epsilon},-a_{ss}\Delta t\left(I_{{\mathcal{R}}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)}{R}_{V}^{\top}\right\|,$
    where $R_{U}$ and $R_{V}$ are upper triangular matrices from a reduced Q-less QR decomposition as described in (24), which can be done once for all DIRK stages.
  5. Step 5.
    Compare the computed residual norm $\|{\bf R}^{(s)}_{\mathscr{L}}\|$ with the given error tolerance $\epsilon_{\texttt{tol}}$. If the tolerance is met, then admit the solution with ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$; otherwise, we go back to Step 1 to further augment the xKrylov subspaces.